A bicycle trip of 90 miles would have taken an hour less if the average speed had been increased by 3mi/h. Find the average speed of the biclycle
Answer by mananth(16949)

speed x
speed x - 3

Distance = same 90 miles
original time – time with increased speed =1
t=d/t
90/x-90 /(x+3) =1
LCD= x(x+3 )
multiply by LCD
90(x+3)-90x=x(x +3)
90x+270-90x=X^2 +3x
270 = 1 x^2 + 3 x
x^2 + 3 x - 270 = 0

a= 1 b= 3 c= -270

x1=(-3+sqrt(9-(-1080))/ 2
= 1089
x1= 15

x2=(3-sqrt(9-(-1080)) / 2

x2= -18
ignore negative
speed = 15 mph
m.ananth@hotmail.ca